repeat 1. Represent 23.689 as a quotient of two integers. (Calculate 2. Find the limit (if it exists) lim (1 + 1)n+1 n +00
Answer #1

© 23.689 = 23.689 689689... a 23+ 680 + 4COX160 + 1.600*0.600) = 23 +680 (1 + 1000 1000 12 + ....) 97 23 + 689 = 23675 0999 2